# Almost device-independent certification of GME states with minimal measurements

### 28 Feb 2024 v1

@@ III.1 Graph states
Figure 2: Examples of graph: (a) ring graph (b) star graph corresponding to $N-$qudit GHZ state up to local unitaries.
Let us begin by defining graph states of arbitrary local dimension [42, 35, 43]. Consider a multigraph $G=(V,E,\Gamma)$ (cf. Fig. 2), where $V$ is a set of all the vertices of the graph, whereas $E$ is a set of pairs of vertices that are connected by edges. Finally, $\Gamma$ is a matrix with elements $\gamma_{i,j}=0,\ldots,d-1$ denoting the multiplicity of the edges between vertices $i$ and $j$. For our purpose, we consider graphs that are connected, that is, those in which there exist of path between any pair of vertices. Then, the graph state $|G\rangle\in(\mathbbm{C}^{d})^{\otimes N}$ corresponding to the multigraph $G=(V,E,\Gamma)$, where $N$ is the number of vertices, is constructed as:
  • Each vertex of $G$ is associated the state $|+^{d}_{i}\rangle=\frac{1}{\sqrt{d}}\sum_{i=0}^{d-1}|i\rangle$ where $\{|i\rangle\}$ is the $d-$dimensional computational basis.
  • For each edge $e={i,j}$ connecting the vertex $i,j$ of multiplicity $\gamma_{i,j}$, a controlled unitary
@@
    is applied on the state
    $|+^{d}_{V}\rangle=\bigotimes_{i=1}^{N}|+_{i}^{d}\rangle.$ (9)
Consequently, a graph state corresponding to the multigraphh $G=(V,E)$ is defined as
$|G\rangle=\prod_{i,j}Z_{i,j}^{\gamma_{i,j}}|+^{d}_{V}\rangle.$ (10)
An equivalent way to represent graph states is by using the stabilizer formalism. Given a multigraph $G=(V,E)$, let us consider a vertex $i$ and associate the operator $X$ to it. All the vertices attached by an edge $e$ to the $i-$th vertex are associated with the operator $Z$. Thus, the stabilizing operator corresponding to the $i-$th vertex is given by
$S_{i}(G)=X_{i}\otimes\bigotimes_{j\in n(i)}Z_{j}^{\gamma_{i,j}}\qquad(i=1,2,\ldots,N),$ (11)
where $n(i)$ is the set of vertices connected to the vertex $i$. The graph state $|G\rangle$ is the unique eigenstate of all the stabilizing operators $S_{i}(G)$ corresponding to the eigenvalue $1$. For example, the well-known Greenberger-Horne-Zeilinger (GHZ) state is a graph state.
Let us now move on to constructing steering inequalities that are maximally violated by any graph state and then use them later for self-testing. For this purpose, we construct the steering operator by considering the stabilizing operators of a graph state of the form (11). Then, we consider the first vertex, and label it as 00 vertex, to be trusted and thus the measurement operators acting on it are known to be $Z_{d}$ and $X_{d}$ (1). Next, we replace $Z_{d}$ and $X_{d}$ on the $j$-th vertex with $B_{j,0},B_{j,1}$ respectively. The steering operator is then given by
$\mathcal{I}_{d}(G,N)$ $=$ $X_{A}\otimes\bigotimes_{j\in n(0)}(B_{j,0})^{\gamma_{0,j}}$ (12)
@@
where $h.c.$ denotes the Hermitian conjugate of the operator on its left-hand side with $G,N$ denoting the graph and number of vertices respectively. Notice that, given a graph $G=(V,E)$, the first term of the above operator corresponds to the stabilizer when choosing the vertex 00, the second term consists of the stabilizers when choosing vertices connected to the $0^{\text{th}}$ vertex and the third term consists of rest of it. Also notice that in the above operator $\gamma_{i,j}=\gamma_{j,i}$ for any $i,j$. Thus, the steering inequality is given by
@@
where $\beta_{L}$ is the maximum value attainable using the LHS model. In Fact 1 of Appendix A, we provide an upper bound to $\beta_{L}$ in terms of a simple optimization problem that can be numerically evaluated for any graph $G$ and any $N,d$. The quantum bound $\beta_{Q}$ of the steering functional $\langle\mathcal{I}_{d}(G,N)\rangle$ is equal to the number of terms in $\mathcal{I}_{d}(G,N)$ (12), that is, $\beta_{Q}=2N$ and can be attained by the graph state $|G\rangle$ and observables $B_{j,0}=Z,B_{j,1}=X$ for any $j$. Importantly, from Fact 1 of Appendix A, we can conclude that $\beta_{L}<\beta_{Q}$ and thus the proposed steering inequality is non-trivial.
@@
$U_{j}\,B_{j,0}\,U_{j}^{\dagger}=Z_{B^{\prime}_{j}}\otimes\mathbbm{1}_{B^{\prime\prime}_{j}},\quad U_{j}\,B_{j,1}\,U_{j}^{\dagger}=X_{B^{\prime}_{j}}\otimes\mathbbm{1}_{B^{\prime\prime}_{j}}$ (14)
$\left(\mathbbm{1}_{AE}\otimes\bigotimes_{i=1}^{N-1}U_{j}\right)|\psi_{N}\rangle=|G\rangle_{A\mathbf{B}^{\prime}}\otimes|\xi\rangle_{\mathbf{B}^{\prime\prime}E},$ (15)
The proof of the above theorem is given in Appendix A.
@@ III.2 Schmidt states
Every pure bipartite state $|\psi\rangle_{AB}$ admits a Schmidt decomposition, that is, there exist orthonormal bases $|i\rangle_{A}$ and $|i\rangle_{B}$ in $\mathcal{H}_{A}$ and $\mathcal{H}_{B}$, respectively, such that $|\psi\rangle_{AB}=\sum_{i}\alpha_{i}|i\rangle_{A}\otimes|i\rangle_{B}$, where $\alpha_{i}>0$ and $\sum_{i}\alpha_{i}^{2}=1$. In the multipartite setting, an analogous decomposition is in general impossible. The multipartite pure states which can be represented via Schmidt decomposition are referred to as Schmidt states [17]. Under the action of local unitary operators, a multipartite Schmidt state $|\psi(\boldsymbol{\alpha})\rangle_{A\mathbf{B}}$ belonging to $\mathbb{C}^{d_{A}}\otimes\mathbb{C}^{d_{B_{1}}}\ldots\otimes\mathbb{C}^{d_{B_{N-1}}}$ can be reduced to the following form
@@
where $d=\min\{d_{A},d_{B_{1}},\ldots,d_{B_{N-1}}\}$ and $\boldsymbol{\alpha}\equiv\alpha_{1},\ldots,\alpha_{N}$, such that $10<\alpha_{i}<$ 1 for all $i$ and $\sum_{i=0}^{d-1}\alpha_{i}^{2}=1$. Because $\alpha_{i}>0$ for all $i$, the Schmidt rank of the state is maximal, which ensures that it is a $N$-qudit state.
Inspired from [31], let us now introduce a general class of steering inequalities that are maximally violated by states of the form (16) for a given $\boldsymbol{\alpha}$. For this purpose, we consider the following steering operator
$\mathcal{S}_{d}(\boldsymbol{\alpha},N)$ $=$ $\sum_{k=1}^{d-1}\left[\left(\sum_{j=1}^{N-1}A_{0}^{k}\otimes B_{j,0}^{k}\right)\right.$ (17)
$+\left.\gamma(\boldsymbol{\alpha})A_{1}^{k}\otimes\bigotimes_{j=1}^{N-1}B_{j,1}^{k}+\delta_{k}(\boldsymbol{\alpha})A_{0}^{k}\right],$
@@
In our scenario, we consider that all the $N-1$ Bobs perform two arbitrary measurements denoted by $B_{i,0}$ and $B_{i,1}$. Also, we assume that the local dimensions of the state shared by them are unknown. At the same time, because Alice’s device is trusted, her measurements are completely characterized and the local dimension of the shared state on Alice’s side is known. Alice’s observables are the generalized Pauli operators (1), that is, $A_{0}=Z_{d}$ and $A_{1}=X_{d}$.
@@
Again $\beta_{L}$ is the maximum value of $\langle\mathcal{S}_{d}(\boldsymbol{\alpha},N)\rangle$ attainable with LHS models. In Fact 3 of Appendix B, we provide a form of $\beta_{L}$ in terms of an optimization problem that can be numerically evaluated for any $d,N,\boldsymbol{\alpha}$. Then, the quantum bound $\beta_{Q}$ of $\langle\mathcal{S}_{d}(\boldsymbol{\alpha},N)\rangle$ is given by $\beta_{Q}=(d-1)(N-1)+1$ and can be attained by $|\psi(\boldsymbol{\alpha})\rangle$ and observables $B_{j,0}=Z_{d}$ and $B_{j,1}=X_{d}$ for any $j$. Importantly, from Fact 3 we can conclude that $\beta_{L}<\beta_{Q}$ and thus the proposed steering inequality is non-trivial.
Now, if all the parties observe that $\langle\mathcal{S}_{d}(\boldsymbol{\alpha},N)\rangle=(d-1)(N-1)+1$, then we can establish the following theorem.
@@
[…]
$U_{i}\,B_{i,0}\,U_{i}^{\dagger}=Z_{d}\otimes\mathbbm{1}_{B^{\prime\prime}_{i}},\quad U_{i}\,B_{i,1}\,U_{i}^{\dagger}=X_{d}\otimes\mathbbm{1}_{B^{\prime\prime}_{i}},$ (25)
$\left(\mathbbm{1}_{AE}\otimes\bigotimes_{i=1}^{N-1}U_{j}\right)|\psi_{N}\rangle=|\psi(\boldsymbol{\alpha})\rangle_{A\mathbf{B}^{\prime}}\otimes|\xi\rangle_{\mathbf{B}^{\prime\prime}E},$ (26)
@@ III.3 $N$-qubit generalized $W$ states
The third and last class of states that we are going to consider comprises the $N$-qubit generalized $W$ states. As the name suggests, they represent a generalization of the well-known $W$ state $|\psi_{W}\rangle=(|100\rangle+|010\rangle+|001\rangle)/\sqrt{3}$ and have the following form
$|\psi_{W_{N}}\rangle=\sum_{l=0}^{N-1}\alpha_{l+1}|0\ldots,0,1_{l},0,\ldots,0\rangle,$ (27)
@@
$\gamma_{l}=\frac{2\alpha_{l+1}\alpha_{1}}{\alpha_{l+1}^{2}+\alpha_{1}^{2}},\quad\delta_{l}=\frac{\alpha_{l+1}^{2}-\alpha_{1}^{2}}{\alpha_{l+1}^{2}+\alpha_{1}^{2}},$ (29)
@@
where $\beta_{L}$ is the LHS bound. Unfortunately, we were not able to calculate the LHS bound for this case. The quantum bound, nonetheless, is given by $\beta_{Q}=2N$ and is achieved by the state $|\psi_{W_{N}}\rangle$ and measurements $B_{j,0}=Z$ and $B_{j,1}=X$, as we prove in Fact 5 of Appendix C. By observing this value, the parties can certify the state and measurements through the following statement:
@@
[…]
$U_{j}\,B_{j,0}\,U_{j}^{\dagger}=Z_{B^{\prime}_{j}}\otimes\mathbbm{1}_{B^{\prime\prime}_{j}},\quad U_{j}\,B_{j,1}\,U_{j}^{\dagger}=X_{B^{\prime}_{j}}\otimes\mathbbm{1}_{B^{\prime\prime}_{j}}$ (32)

### 13 Mar 2026 v2

@@ III.1 Graph states
Figure 2: Examples of graph: (a) ring graph (b) star graph corresponding to NN-qudit GHZ state up to local unitaries.
Let us begin by defining graph states of arbitrary local dimension [47, 48, 49]. Consider a multigraph $G=(V,E,\Gamma)$ (cf. Fig. 2), where $V$ is a set of all the vertices of the graph, whereas $E$ is a set of pairs of vertices that are connected by edges. Finally, $\Gamma$ is a matrix with elements $\gamma_{i,j}\in\{0,\ldots,d-1\}$ specifying the multiplicities of the edges between vertices $i$ and $j$; notice that $\gamma_{i,j}=\gamma_{j,i}$ for any pair $i,j$ and $\gamma_{i,i}=0$. For our purpose, we consider graphs that are connected, that is, those in which there exists a path between any pair of vertices. Then, the graph state $|G\rangle\in(\mathbbm{C}^{d})^{\otimes N}$ corresponding to the multigraph $G=(V,E,\Gamma)$, where $N$ is the number of vertices, is constructed as:
  • Each vertex of $G$ is associated with a single qudit state $|+^{d}_{i}\rangle=(1/\sqrt{d})\sum_{i=0}^{d-1}|i\rangle$, where $\{|i\rangle\}$ is the $d$-dimensional computational basis. The todal state of $N$ qudits is thus
    $|+^{d}_{V}\rangle=\bigotimes_{i=1}^{N}|+_{i}^{d}\rangle.$ (10)
  • For each edge $e=\{i,j\}$ connecting the vertices $i$ and $j$ of multiplicity $\gamma_{i,j}$, one applies the controlled unitary
@@
    is applied to the state $|+^{d}_{V}\rangle$.
Consequently, a graph state corresponding to the multigraph $G=(V,E)$ is defined as
$|G\rangle=\prod_{\{i,j\}\in E}Z_{i,j}^{\gamma_{i,j}}|+^{d}_{V}\rangle.$ (12)
An equivalent way to represent graph states is by using the stabilizer formalism [50]. Let us consider a multigraph $G=(V,E)$ and associate to each vertex $i$ the following operator
$\overline{S}(G)_{d,i}=X_{i}\otimes\bigotimes_{j\in n(i)}Z_{j}^{\gamma_{i,j}}\qquad(i=1,2,\ldots,N),$ (13)
where $n(i)$ is the set of vertices connected to the vertex $i$, which is also called the neighborhood of $i$. The graph state $|G\rangle$ is the unique eigenstate of all the $N$ stabilizing operators $S_{i}(G)$ corresponding to the eigenvalue $1$. For example, the GHZ states are graph states.
Before presenting our first results, we note that fully device-independent schemes for graph states of local dimension two or three have already been proposed in Refs. [15, 14] and [19], respectively, whereas no scheme is known beyond these specific dimensions. Here, we provide strategies that allow for the certification of graph states of arbitrary local dimension, albeit under the assumption that one of the measurement devices is trusted.
Let us now move on to constructing steering inequalities that are maximally violated by graph states and then use these inequalities for self-testing. Let us consider a graph state $|G\rangle$ associated to a graph $G$. We construct the steering operator from the corresponding stabilizing operators (13). We then assume the first vertex to be trusted and label it as 0; thus the measurement operators acting on the first site are known to be $Z_{d}$ and $X_{d}$ (1). Next, we replace $Z_{d}$ and $X_{d}$ on the $j$-th vertex with arbitrary $d$-outcome observables $B_{j,0}$ and $B_{j,1}$, respectively. The steering operator is then given by
$\mathcal{I}_{d}(G,N)$ $=$ $X_{A}\otimes\bigotimes_{j\in n(0)}(B_{j,0})^{\gamma_{0,j}}$
@@
where $h.c.$ denotes the Hermitian conjugate of the operator on its left-hand side. Notice that for a given graph $G=(V,E)$, the first term of the above operator corresponds to the stabilizer when choosing the vertex 0, the second term consists of the stabilizers when choosing vertices connected to the 0th vertex and the third term consists of rest of it. Also notice that in the above operator $\gamma_{i,j}=\gamma_{j,i}$ for any $i,j$. Thus, the steering inequality is given by
@@
where $\beta_{L}$ is the maximum value attainable using the LHS model. In Fact 1 of Appendix A, we provide an upper bound to $\beta_{L}$ in terms of a simple optimization problem that can be numerically evaluated for any graph $G$ and any number of parties $N$ and outcomes $d$.
The quantum bound $\beta_{Q}$ of the steering functional $\langle\mathcal{I}_{d}(G,N)\rangle$ is equal to the number of terms it consists of, which is twice the number of the corresponding stabilizing operators, that is, $\beta_{Q}=2N$. Clearly, the quantum bound is attained by the graph state $|G\rangle$ and observables $B_{j,0}=Z$ and $B_{j,1}=X$ for any $j$. Importantly, it follows from Fact 1 of Appendix A that $\beta_{L}<\beta_{Q}$, and consequently the proposed steering inequality is non-trivial for any graph state.
@@
Assume that for a given graph $G$, the corresponding functional (III.1) reaches its maximal value $\langle\mathcal{I}_{d}(G,N)\rangle=2N$ for a state $|\psi_{A\mathbf{B}E}\rangle$ and observables $B_{i,y_{i}}$. Then, the following statement holds true for any $d$: $\mathcal{H}_{B_{i}}=(\mathbbm{C}^{d})_{B^{\prime}_{i}}\otimes\mathcal{H}_{B^{\prime\prime}_{i}}$, where $\mathcal{H}_{B^{\prime\prime}_{i}}$ is some finite-dimensional Hilbert space, and there exists a local unitary transformation on Bob’s side $U_{i}:\mathcal{H}_{B_{i}}\rightarrow\mathcal{H}_{B_{i}}$, such that for any $i$,
$U_{i}\,B_{i,0}\,U_{i}^{\dagger}=Z\otimes\mathbbm{1}_{B^{\prime\prime}_{i}},\quad U_{i}\,B_{i,1}\,U_{i}^{\dagger}=X\otimes\mathbbm{1}_{B^{\prime\prime}_{i}},\ \$(16)
where $B^{\prime\prime}$} denotes Bob’s auxiliary system and the state $|\psi\rangle_{A\mathbf{B}E}$ is given by,
$\left(\mathbbm{1}_{AE}\otimes\bigotimes_{i=1}^{N-1}U_{i}\right)|\psi\rangle_{A\mathbf{B}E}=|G\rangle_{A\mathbf{B}^{\prime}}\otimes|\xi\rangle_{\mathbf{B}^{\prime\prime}E},$ (17)
where $|\xi\rangle_{\mathbf{B}^{\prime\prime}E}$ denotes the auxiliary state corresponding to subsystems $\mathbf{B}^{\prime\prime}$ and $E$.
The proof of the above theorem is given in Appendix A.
In practical scenarios, it is important to investigate the robustness of self-testing statements, such as those introduced in our work, against experimental imperfections, including noise in the preparation device and imperfect measurements. However, deriving analytical robustness bounds is typically a highly demanding task, especially for systems of arbitrary local dimension. In the simplest case of graph states of local dimension two, we can instead exploit Jordan’s lemma, which significantly simplifies the problem and allows us to obtain the following robustness bound.
Theorem 2.
Assume that the steering inequality (41) attains a value close to the maximal violation, $\langle\mathcal{I}_{2}(G,N)\rangle\geq\beta_{Q}-\varepsilon$, when the $0^{\text{th}}$ party Alice measures the observables $A_{0}=Z$ and $A_{1}=X$. Then, there exist local unitary operations $U_{i}$ such that the following inequality is satisfied,
@@
where, as before,
$|\psi^{\prime}\rangle_{A\mathbf{B}E}=\mathbbm{1}_{AE}\otimes\bigotimes_{i=1}^{N-1}U_{i}|\psi\rangle_{A\mathbf{B}E},$ (19)
$|G\rangle_{A\mathbf{B}^{\prime}}$ is the graph state to be certified and $|\xi\rangle_{\mathbf{B}^{\prime\prime}E}$ is the auxiliary state. Finally $g$ is a factor depending on the particular graph (c.f. Theorem 2 in Appendix A).
A detailed version of the theorem and the proof is given in Appendix A.
@@ III.2 Schmidt states
Every pure bipartite state $|\psi\rangle_{AB}$ admits a Schmidt decomposition, that is, there exist orthonormal bases $|i\rangle_{A}$ and $|i\rangle_{B}$ in $\mathcal{H}_{A}$ and $\mathcal{H}_{B}$, respectively, such that $|\psi\rangle_{AB}=\sum_{i}\alpha_{i}|i\rangle_{A}\otimes|i\rangle_{B}$, where $\alpha_{i}>0$ and $\sum_{i}\alpha_{i}^{2}=1$. An analogous decomposition in the multipartite setting is in general impossible. The multipartite pure states which can be represented via Schmidt decomposition are referred to as Schmidt states [17]. Under the action of local unitary operators, an $N$-qudit Schmidt state $|\psi(\boldsymbol{\alpha})\rangle_{A\mathbf{B}}$ can be brought to the following form
@@
$\boldsymbol{\alpha}\equiv\alpha_{1},\ldots,\alpha_{N}$ such that $0<\alpha_{i}<1$ for all $i$ and $\sum_{i=0}^{d-1}\alpha_{i}^{2}=1$. Because $\alpha_{i}>0$ for all $i$, the Schmidt rank of the state is maximal, which ensures that it is a truly $N$-qudit state.
Importantly, a self-testing scheme for multipartite Schmidt states in the device-independent regime is already known and was proposed in Ref. [17]. That approach, however, requires more than two measurements per observer and is therefore suboptimal from an implementation perspective. Specifically, N1N-1 parties perform three measurements each, while the remaining party performs four. In this work, we exploit the assumption that one of the observers is trusted to reduce the number of measurements per observer to two.
Inspired by Ref. [31], let us now introduce a general class of steering inequalities that are maximally violated by the above Schmidt state for a given $\boldsymbol{\alpha}$. For this purpose, we consider the following steering operator
$\mathcal{S}_{d}(\boldsymbol{\alpha},N)$ $=$ $\sum_{k=1}^{d-1}\left[\left(\sum_{i=1}^{N-1}A_{0}^{k}\otimes B_{i,0}^{k}\right)\right$. (21)
$+\left.\gamma(\boldsymbol{\alpha})A_{1}^{k}\otimes\bigotimes_{i=1}^{N-1}B_{i,1}^{k}+\delta_{k}(\boldsymbol{\alpha})A_{0}^{k}\right],$
@@
In our scenario, we consider that all Bobs perform two arbitrary measurements, denoted by $B_{i,0}$ and $B_{i,1}$, whereas Alice’s measurement device is trusted and measures the generalized Pauli observables $A_{0}=Z$ and $A_{1}=X$.
@@
Whereas it is a difficult task to analytically compute the maximal classical value of the expression $\beta_{L}$, in Fact 3 of Appendix B we provide an upper bound on $\beta_{L}$ in terms of an optimization problem that can be numerically evaluated. Then, the quantum bound $\beta_{Q}$ of $\langle\mathcal{S}_{d}(\boldsymbol{\alpha},N)\rangle$ is given by $\beta_{Q}=(d-1)(N-1)+1$ and can be attained by $|\psi(\boldsymbol{\alpha})\rangle$ and observables $B_{j,0}=Z_{d}$ and $B_{j,1}=X_{d}$ for any $j$. Importantly, from Fact 3 we can also conclude that $\beta_{L}<\beta_{Q}$ and thus the proposed steering inequality is non-trivial.
Let us now state our result.
@@
Assume that the steering inequality (24) is maximally violated, that is, $\langle\mathcal{S}_{d}(\boldsymbol{\alpha},N)\rangle=(d-1)(N-1)+1$. Then, the following statement holds true for any $d$: $\mathcal{H}_{B_{i}}=(\mathbbm{C}^{d})_{B^{\prime}_{i}}\otimes\mathcal{H}_{B^{\prime\prime}_{i}}$, where $\mathcal{H}_{B^{\prime\prime}_{i}}$ is a finite-dimensional Hilbert space, and there exists a local unitary transformation on Bob’s side $U_{i}:\mathcal{H}_{B_{i}}\rightarrow\mathbbm{C}^{d}\otimes\mathcal{H}_{B^{\prime\prime}_{i}}$ such that for any $i$,
$U_{i}\,B_{i,0}\,U_{i}^{\dagger}=Z\otimes\mathbbm{1}_{B^{\prime\prime}_{i}},\quad U_{i}\,B_{i,1}\,U_{i}^{\dagger}=X\otimes\mathbbm{1}_{B^{\prime\prime}_{i}},$ (25)
where $B^{\prime\prime}_{i}$ denotes Bob’s auxiliary system and the state $|\psi\rangle_{A\mathbf{B}}$ is given by
$\left(\mathbbm{1}_{AE}\otimes\bigotimes_{i=1}^{N-1}U_{i}\right)|\psi_{N}\rangle=|\psi(\boldsymbol{\alpha})\rangle_{A\mathbf{B}^{\prime}}\otimes|\xi\rangle_{\mathbf{B}^{\prime\prime}E},$ (26)
where $|\xi\rangle_{\mathbf{B}^{\prime\prime}E}$ denotes the auxiliary state.
@@ III.3 $N$-qubit generalized $W$ states
The third and last class of states that we consider comprises the $N$-qubit generalized $W$ states. As the name suggests, they represent a generalization of the well-known $W$ state $|\psi_{W}\rangle=(|100\rangle+|010\rangle+|001\rangle)/\sqrt{3}$ and have the following form
$|\psi_{W}(\boldsymbol{\alpha})\rangle=\sum_{l=0}^{N-1}\alpha_{l+1}|0\ldots,0,1_{l},0,\ldots,0\rangle,$ (27)
@@
$\gamma_{l}=\frac{2\alpha_{l+1}\alpha_{1}}{\alpha_{l+1}^{2}+\alpha_{1}^{2}},\qquad\delta_{l}=\frac{\alpha_{l+1}^{2}-\alpha_{1}^{2}}{\alpha_{l+1}^{2}+\alpha_{1}^{2}},$ (29)
@@
and $B_{k,0}$ and $B_{k,1}$ are the measurements acting on the $k$-th qubit. Similarly, $\mathbbm{1}_{l}$ is the identity operator acting on the $l$-th qubit. The corresponding steering inequality is given by
@@
While as before we are unable to compute the LHS bound $\beta_{L}$, we can prove the inequality to be nontrivial by demonstrating that observation of the maximal quantum value of $\langle\mathcal{W}_{N}(\boldsymbol{\alpha})\rangle$, which is $\beta_{Q}=2N$ (see Appendix C for a proof), allows one to self-test $|\psi_{W}(\boldsymbol{\alpha})\rangle$. In fact, we can prove the following statement.
@@
Assume that $\langle\mathcal{W}_{N}(\boldsymbol{\alpha})\rangle=2N$ for a state $|\psi\rangle_{A\mathbf{B}E}$ and Bob’s observables $B_{i,y_{i}}$. Then, $\mathcal{H}_{B_{i}}=\mathbbm{C}^{2}\otimes\mathcal{H}_{B^{\prime\prime}_{i}}$, where $\mathcal{H}_{B^{\prime\prime}_{i}}$ is some finite-dimensional Hilbert space, and there exists a local unitary transformation on Bob’s side $U_{i}:\mathcal{H}_{B_{i}}\rightarrow\mathbbm{C}^{2}\otimes\mathcal{H}_{B^{\prime\prime}_{i}}$, such that
$U_{i}\,B_{i,0}\,U_{i}^{\dagger}=Z\otimes\mathbbm{1}_{B^{\prime\prime}_{i}},\quad U_{i}\,B_{i,1}\,U_{i}^{\dagger}=X\otimes\mathbbm{1}_{B^{\prime\prime}_{i}}$ (32)